# Cultivating $|T\rangle$ states on the surface code with only two-qubit gates

### v1

@@ I Introduction
Two recent proposals have made partial progress in surface code cultivation. Both of these proposals are tailored to neutral atoms, requiring the flexible connectivity enabled by reconfigurable qubits. First, Ref. [20] proposed a method for cultivating two-qubit $|CX\rangle$ magic states on two patches of surface code. However, this method has three drawbacks: (1) it uses physical three-qubit CCX gates, which are more difficult to implement than two-qubit gates; (2) $|CX\rangle$ states cannot be used directly to teleport gates, but must be probabilistically converted into Toffoli states [21, 22]; (3) it produces $|CX\rangle$ states that are roughly $10^{3}\times$ noisier than the $|T\rangle$ states produced by color code cultivation. Second, Ref. [2] proposed a method for cultivating $|T\rangle$ states on a variant of the surface code using only two-qubit gates; this method achieves similar fidelities to color-code cultivation and had an overall lower spacetime volume per $|T\rangle$ state, demonstrating the promise of surface code cultivation over color code cultivation. However, this method performs cultivation on the surface code defined on a non-orientable surface (the real projective plane, $\mathbb{RP}^{2}$) before escaping to the usual planar surface code. One significant drawback in this approach is that the non-orientable boundary conditions of $\mathbb{RP}^{2}$ mean the cultivation circuit requires many long-distance connections that are difficult to implement, even in neutral atoms. In addition, low-depth unitary injection and expansion circuits are not known for the $\mathbb{RP}^{2}$ surface code, so that state injection and intermediate code growth must be done with measurement-based circuits which require mid-circuit measurements and feedback, significantly slowing the procedure in neutral atom platforms with slow measurements.
@@ II Results
Second, we note that both the uniform depolarizing and the depolarizing no-idle error models likely do not capture the relevant features of neutral atom qubits. A more accurate error model would include noise induced by qubit movement and transfers between traps, in which the relative performance of the $\mathbb{RP}^{2}$ circuits would likely degrade but the performance of the color circuits may improve somewhat. These simulations also do not account for qubit leakage, which becomes increasingly damaging as the circuit depth increases. On the other hand, we also do not currently take advantage of the structure of noise available in certain in neutral atoms, such as erasure bias [28, 29, 30, senoo2025high, muniz2025high], which may improve performance [31, 38]. As a rough approximation, if a fraction $R_{e}$ of errors are erasures, the leading order error rate of the ungrown patch will go from $Cp^{d}$ to $C((1-R_{e})p)^{d}$ for some constant $C$, meaning that a realistic $90\%$ erasure fraction [28] could improve the logical error rate of $d=5$ cultivation by a factor of $10^{5}$ (see also [38]).
@@ III.1 Injection and code growth
Figure 4: Unitary circuits to grow the surface code from $d=3$ to $d=4$ (top) or $d=5$ (bottom).
To grow a $d=3$ surface code to a $d=4$ or $d=5$ surface code, we use the unitary growth circuits illustrated in Fig. 4, which are similar to the growth circuits in Ref. [33] but lower depth. Unitary growth also avoids the midcircuit measurement and feedback used to grow the $d=3$ $\mathbb{RP}^{2}$ code to a $d=5$ $\mathbb{RP}^{2}$ code in Ref. [2].
@@ IV Conclusion
We also re-emphasize that beyond the noise model model, the measure of spacetime volume in previous works is ad-hoc and not necessarily relevant to neutral atom qubit platforms, More accurate spacetime comparisons will require careful consideration of atom move times and measurement times.
@@
While completing this manuscript, we became aware of two pieces of related work by the Puri and AWS research groups, which appear in the same arXiv posting as the present work.

### v2

@@ I Introduction
Two recent proposals have made partial progress in surface code cultivation. Both of these proposals are tailored to neutral atoms, requiring the flexible connectivity enabled by reconfigurable qubits. First, Ref. [20] proposed a method for cultivating two-qubit $|CX\rangle$ magic states on two patches of surface code. However, this method has three drawbacks: (1) it uses physical three-qubit CCX gates, which are more difficult to implement than two-qubit gates; (2) $|CX\rangle$ states cannot be used directly to teleport gates, but must be probabilistically converted into Toffoli states [21, 22]; (3) it produces $|CX\rangle$ states that are roughly $10^{3}\times$ noisier than the $|T\rangle$ states produced by color code cultivation. Second, Ref. [2] proposed a method for cultivating $|T\rangle$ states on a variant of the surface code using only two-qubit gates; this method achieves similar fidelities to color-code cultivation and had an overall lower spacetime volume per $|T\rangle$ state, demonstrating the promise of surface code cultivation over color code cultivation. However, this method performs cultivation on the surface code defined on a non-orientable surface (the real projective plane, $\mathbb{RP}^{2}$, see also [23]) before escaping to the usual planar surface code. One significant drawback in this approach is that the non-orientable boundary conditions of $\mathbb{RP}^{2}$ mean the cultivation circuit requires many long-distance connections that are difficult to implement, even in neutral atoms. In addition, low-depth unitary injection and expansion circuits are not known for the $\mathbb{RP}^{2}$ surface code, so that state injection and intermediate code growth must be done with measurement-based circuits which require mid-circuit measurements and feedback, significantly slowing the procedure in neutral atom platforms with slow measurements.
@@ II Results
Second, we note that both the uniform depolarizing and the depolarizing no-idle error models likely do not capture the relevant features of neutral atom qubits. A more accurate error model would include noise induced by qubit movement and transfers between traps, in which the relative performance of the $\mathbb{RP}^{2}$ circuits would likely degrade but the performance of the color circuits may improve somewhat. These simulations also do not account for qubit leakage, which becomes increasingly damaging as the circuit depth increases. On the other hand, we also do not currently take advantage of the structure of noise available in certain in neutral atoms, such as erasure bias [30, 31, 32, 33, 34], which may improve performance [35, 36]. As a rough approximation, if a fraction $R_{e}$ of errors are erasures, the leading order error rate of the ungrown patch will go from $Cp^{d}$ to $C((1-R_{e})p)^{d}$ for some constant $C$, meaning that a realistic $90\%$ erasure fraction [30] could improve the logical error rate of $d=5$ cultivation by a factor of $10^{5}$ (see also [36]).
@@ III.1 Injection and code growth
Figure 4: Unitary circuits to grow the surface code from $d=3$ to $d=4$ (top) or $d=5$ (bottom). Modified versions of the low-depth injection circuits introduced in [38].
To grow a $d=3$ surface code to a $d=4$ or $d=5$ surface code, we use the unitary growth circuits illustrated in Fig. 4 which are modified versions of the low-depth circuits introduced in [38]. Unitary growth also avoids the midcircuit measurement and feedback used to grow the $d=3$ $\mathbb{RP}^{2}$ code to a $d=5$ $\mathbb{RP}^{2}$ code in Ref. [2].
@@ IV Conclusion
We also re-emphasize that beyond the noise model, the measure of spacetime volume in previous works is ad-hoc and not necessarily relevant to neutral atom qubit platforms, More accurate spacetime comparisons will require careful consideration of atom move times and measurement times.
@@
While completing this manuscript, we became aware of two pieces of related work by the Puri group [46] and researchers at AWS [updated version of 20], which appeared in the same arXiv posting as the present work.